Alvin is 5 years older than Joseph. In 15 years the sum of their ages is 65. Find their present ages
Answer by Cromlix(4381)

Alvin is 5yrs older than Joseph
A = J + 5
In 15yrs the sum of their ages will be 65
A + 15 + J + 15 = 65
Replace A with J + 5 from
first equation.
J + 5 + 15 + J + 15 = 65
Collect like terms.
J + J = 65 - 5 - 15 - 15
2J = 30
J = 15
Joseph = 15yrs old
Alvin is 20yrs old.
Hope this helps.
:-)